# End-to-end Learnable Clustering for Intent Learning in Recommendation

### 11 Jan 2024 v1

@@ 5.1. Industrial Result
We test the online recommendation performance of our proposed module. Concretely, the overall recommendation performance is listed in Table 4. The metrics contain two parts, including live-streaming metrics (PVCTR and VV) and commercial metrics (PVCTR and UVCTR). “B” denotes the baseline and “B+ELCM” denotes the baseline with the proposed end-to-end learnable cluster module. The results indicate that our proposed module can improve the recommendation performance of the baseline by about 2%. Besides, “B+ELCM” outperforms “B” significantly with $p<0.05$ with three out of four metrics.
@@ 6. Conclusion
The topic of this paper is intent learning in recommendation systems. Concretely, the intent learning methods aim to mine the users’ underlying intents and perform recommendations based on their learned intents. In this paper, we carefully analyze the drawbacks of previous methods. The complex and cumbersome alternating optimization limits the scalability and performance of the existing method. To this end, we propose an end-to-end learnable clustering framework named ELCRec for intent learning in recommendation systems. To be specific, the cluster centers are initialized as the learnable neural parameters. Then, the proposed clustering loss optimizes the cluster distribution based on the mini-batch data, enabling the scalability of ELCRec. Besides, to further improve the recommendation performance, the learned cluster centers are assigned as the self-supervision signals for sequential representation learning. Extensive experiments demonstrate ELCRec’s six abilities, including superiority, effectiveness, efficiency, sensitivity, and convergence. In addition, benefiting from the versatility of ELCRec, we successfully apply it to the real-world industrial scenario and also achieve promising performance. In the future, it is worth developing the cluster number-free clustering framework for intent learning based on the sample density.
@@ A.1. Notation & Dataset
$\mathcal{L}_{\text{overall}}$ Overall loss of the proposed ELCRec
@@
$\mathbf{C}\in\mathbb{R}^{k\times d}$ Learnable cluster center embeddings
@@ A.2. Algorithm
Due to the limitation of the pages, we present the overall algorithm of the proposed ELCRec method in Algorithm 1.
Algorithm 1 End-to-end Learnable Clustering Framework for Recommendation (ELCRec)
@@
9:     # End-to-end Learnable Cluster Module
@@
23:  Return Trained ELCRec model.
@@ A.3. Deployment Details
We aim to apply our proposed method to the real-world industrial scenario. Concretely, the ELCRec algorithm is applied to the front page of the Alipay app. The user view (UV) and page view (PV) of this application are about 50 million and 130 million, respectively. Since most of the users are new to this application, it easily leads to the sparsity of users’ behaviors, namely, the cold-start problem in recommendation systems. Our proposed ELCRec can alleviate this problem by grouping users and then making recommendations. Firstly, we introduce the online baseline of this project. Since the sparsity of the users’ behaviors, we replaced the users’ behaviors with the users’ activities. Then, the online baseline multi-gate mixture-of-expert (MMOE) (MMOE, ) models the users’ activities. In this model, the experts are designed to extract the features of users, and the multi-gates are designed to select specific experts. The inputs of the multi-gates are the activities of the users. This design aims to train an activity-awarded model to group different users and then conduct recommendations. However, we found the performance of this model is limited, and the output of the gates is smooth, indicating that this model may fail to group users. Meanwhile, on the open benchmarks, extensive experiments demonstrate the proposed end-to-end learnable clustering module is effective and scalable. Thus, to solve the above problem, ELCRec is adopted in this project. It is designed to assist the gate to group users. For example, the high-activity users and new users are grouped into different clusters, and then the users in different clusters will be recommended differently. Therefore, it alleviates the cold-start issue and further improves the recommendation performance. It is worth mentioning that the networks are trained with multi-task targets, e.g., CTR, CVR, etc. Following the previous online baseline, the method is implemented with the TensorFlow deep learning platform (tensorflow, ).
@@ A.4. Additional Test on Diverse Domains
We acknowledge that the applied domains of our proposed method are limited. We aim to test our proposed ELCRec on other domains, such as news recommendation and movie recommendation.
@@ A.4.1. Movie Recommendation
Specifically, for movie recommendation, we conducted experiments on the MovieLens 1M dataset (ML-1M). This dataset contains 1 million ratings from 6000 users on 4000 movies. In this experiment, we compared our proposed ELCRec with the baseline ICLRec. The experimental results are presented in the table below.
@@
From these experimental results, we can draw three conclusions. (1) Our proposed ELCRec achieves better recommendation performance, as evidenced by higher values for all four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. For example, with the HR@5 metric, ELCRec outperforms ICLRec by 13.56%. (2) We calculated the $p$-value between our method and the baseline ICLRec. The results indicate that all the $p$-values are less than 0.05, suggesting that our ELCRec significantly outperforms ICLRec.
@@ A.5. Sensitivity
This section aims to answer research question (4). To verify the sensitivity of the proposed ELCRec, we test the performance on four datasets with different hyper-parameters. The experimental results are demonstrated in Figure 5. The x-axis denotes the values of hyper-parameters, and the y-axis denotes the values of the HR@5 metric. Concretely, for the trade hyper-parameter $\alpha$, we test the performance with $\alpha\in\{0.01,0.1,1,10,100\}$. We find that our proposed ELCRec is not very sensitive to trade-off $\alpha$. And ELCRec can achieve promising performance when $\alpha\in[0.1,10]$. Besides, for the cluster number $k$, we test the recommendation performance with $\alpha\in\{32,64,128,256,512\}$. The results show that ELCRec is also not very sensitive to cluster number $k$ and can perform well when $k\in[256,512]$.
@@
Figure 5. Sensitivity analyses of ELCRec. The results are evaluated by the HR@5 metric.
@@ A.7. Visualization
To further demonstrate ELCRec’s capability to capture users’ underlying intents, we conduct visualization experiments on four public datasets. Concretely, the learned behavior sequence embeddings are visualized via $t$-SNE during the training process. As shown in Figure 7, the first row to the fourth row denotes the results on Sports, Beauty, Toys, and Yelp, respectively. From these experimental results, we have three observations as follows. (1) At the beginning of training, the behavior sequence embeddings are disorganized and can not reveal the underlying intent of users. (2) During the training process, the latent distribution is optimized, and similar behaviors are grouped into the underlying intent. (3) After training, the users’ underlying intents appear, and we highlight them with circles in Figure 7. These intents can assist recommendation systems in better modeling users’ behavior and recommending items. In summary, the above experiments and observations verify the effectiveness of our proposed ELCRec.
@@ A.9. Solutions to Balance Problem
Balancing the different loss functions in our method is indeed an important challenge. Our overall loss function consists of next-item prediction loss, cluster-assisted contrastive loss, and cluster loss. It is formulated as follows: $\mathcal{L}_{\text{overall}}=\mathcal{L}_{\text{next\_item}}+0.1\times\mathcal{L}_{\text{ccl}}+\alpha\times\mathcal{L}_{\text{cluster}}$. We set the weight of $\mathcal{L}_{\text{ccl}}$ as 0.1 to maintain it in the same order of magnitude as the first term. This reduces the number of hyper-parameters and simplifies the selection process. The weight of $\mathcal{L}_{\text{cluster}}$ is set as a hyper-parameter $\alpha$. We test different values of $\alpha\in\{0.01,0.1,1,10,100\}$ and find that our ELCRec method is not very sensitive to the trade-off $\alpha$. Promising performance is achieved when $\alpha\in[0.1,10]$. The sensitivity analysis experiments are presented in Figure 5 (b) of the Appendix. In our proposed model, we set $\alpha$ to 1 for Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. The selection of $\alpha$ is mainly based on the model performance. Balancing trade-offs in models with multiple loss functions or tasks is a challenge in many deep learning methods. There are various strategies to address this issue.
@@ A.11. Limitations & Future Work
In this paper, we propose a novel intent learning method named ELCRec based on the end-to-end learnable clustering framework. It can better mine users’ underlying intents via unifying representation learning and clustering optimization. Besides, the end-to-end learnable clustering module optimizes the clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment. Moreover, we demonstrate the superiority, effectiveness, efficiency, sensitivity, convergence, and visualization of ELCRec on four benchmarks. ELCRec is also successfully applied in the real-world industrial application.
Although achieving promising results, we admit the proposed ELCRec algorithm has several limitations and drawbacks. We summarize them as follows.
  • Pre-defined Cluster Number. The cluster number in the ELCRec method is a pre-defined hyper-parameter. In the real-world data, it is hard to determine the cluster number, especially under the unsupervised conditions. In this paper, for the open benchmarks, we search the cluster number in 32, 64, 128, 256, 512. Besides, for the real-world application, the cluster number is set to 20 based on the number of user groups. However, either the search method or the expert knowledge can not determine the cluster number well at once. The cluster number may change dynamically during model training, and the proposed method may fail to achieve promising performance.
  • Limited Recommendation Domains. In this paper, we adopt four recommendation datasets, including Sports, Beauty, Toys, and Yelp. But, these four datasets are all business recommendation datasets. Besides, in the industrial application aspect, the used data is also a business recommendation dataset. Therefore, the recommendation domain is limited and this paper merely demonstrates the superiority of ELCRec in one domain.
  • No Strategy for Solving Rapid Preference Change. In real-world applications, the intents of users usually change rapidly. Although the proposed ELCRec can capture the user’s intent effectively in the training data, the user’s intent may change dynamically in the future. However, this work does not provide any strategies to alleviate this issue, which may lead to unpromising performance in real-world applications.
@@
  • More Recommendation Domains. As mentioned above, we admit applied domains of our proposed method are limited. We aim to test our proposed ELCRec on other domains like news recommendation and movie recommendation in the future. For the news recommendation, we can conduct experiments on MIND dataset. This dataset is a large-scale dataset for news recommendation research. It was collected from anonymized behavior logs of the Microsoft News website. MIND contains about 160k English news articles and more than 15 million impression logs generated by 1 million users. For the movie recommendation, we conduct experiments on the MovieLens 1M dataset (ML-1M), Netflix, Douban, Twitch-100k, Twith-full datasets, etc.
  • More Recommendation Domains. As mentioned above, in the real-world application, the intents of the users may change rapidly. However, there is no strategy in the proposed ELCRec method to solve this problem. To this end, in the future, we can propose new cluster center update strategies, such as momentum updates, to control the change rate of the users’ intents. Namely, $\mathbf{C}_{t}=m\cdot\mathbf{C}_{t}+(1-m)\cdot\mathbf{C}_{t-1}$. Here, $\mathbf{C}_{t}$ denote the cluster center embeddings at $t$ and $m$ denotes the momentum. Then, the cluster centers (intents of users) will be changed rapidly when $m$ is large, and the cluster centers (intents of users) will be changed slowly when $m$ is small. This strategy will control the change rate of the users’ intent embeddings, therefore alleviating the above problem.

### 24 Jan 2024 v2

@@ 5.1. Industrial Result
We test the online recommendation performance of our proposed module. Concretely, the overall recommendation performance is listed in Table 4. The metrics contain two parts, including live-streaming metrics (PVCTR and VV) and commercial metrics (PVCTR and UVCTR). “B” denotes the baseline and “B+ELCM” denotes the baseline with the proposed online differentiable cluster module. The results indicate that our proposed module can improve the recommendation performance of the baseline by about 2%. Besides, “B+ELCM” outperforms “B” significantly with $p<0.05$ with three out of four metrics.
@@ 6. Conclusion
The topic of this paper is intent learning in recommendation systems. Concretely, the intent learning methods aim to mine the users’ underlying intents and perform recommendations based on their learned intents. In this paper, we carefully analyze the drawbacks of previous methods. The complex and cumbersome alternating optimization limits the scalability and performance of the existing method. To this end, we propose an online differentiable clustering framework named ODCRec for intent learning in recommendation systems. To be specific, the cluster centers are initialized as the differentiable neural parameters. Then, the proposed clustering loss optimizes the cluster distribution based on the mini-batch data, enabling the scalability of ODCRec. Besides, to further improve the recommendation performance, the learned cluster centers are assigned as the self-supervision signals for sequential representation learning. Extensive experiments demonstrate ODCRec’s six abilities, including superiority, effectiveness, efficiency, sensitivity, and convergence. In addition, benefiting from the versatility of ODCRec, we successfully apply it to the real-world industrial scenario and also achieve promising performance. In the future, it is worth developing the cluster number-free clustering framework for intent learning based on the sample density.
@@ A.1. Notation & Dataset
$\mathcal{L}_{\text{overall}}$ Overall loss of the proposed ODCRec
@@
$\mathbf{C}\in\mathbb{R}^{k\times d}$ Differentiable cluster center embeddings
@@ A.2. Algorithm
Due to the limitation of the pages, we present the overall algorithm of the proposed ODCRec method in Algorithm 1.
Algorithm 1 Online Differentiable Clustering Framework for Recommendation (ODCRec)
@@
9:     # Online Differentiable Cluster Module
@@
23:  Return Trained ODCRec model.
@@ A.3. Deployment Details
We aim to apply our proposed method to the real-world industrial scenario. Concretely, the ODCRec algorithm is applied to the front page of the Alipay app. The user view (UV) and page view (PV) of this application are about 50 million and 130 million, respectively. Since most of the users are new to this application, it easily leads to the sparsity of users’ behaviors, namely, the cold-start problem in recommendation systems. Our proposed ODCRec can alleviate this problem by grouping users and then making recommendations. Firstly, we introduce the online baseline of this project. Since the sparsity of the users’ behaviors, we replaced the users’ behaviors with the users’ activities. Then, the online baseline multi-gate mixture-of-expert (MMOE) (MMOE, ) models the users’ activities. In this model, the experts are designed to extract the features of users, and the multi-gates are designed to select specific experts. The inputs of the multi-gates are the activities of the users. This design aims to train an activity-awarded model to group different users and then conduct recommendations. However, we found the performance of this model is limited, and the output of the gates is smooth, indicating that this model may fail to group users. Meanwhile, on the open benchmarks, extensive experiments demonstrate the proposed online differentiable clustering module is effective and scalable. Thus, to solve the above problem, ODCRec is adopted in this project. It is designed to assist the gate to group users. For example, the high-activity users and new users are grouped into different clusters, and then the users in different clusters will be recommended differently. Therefore, it alleviates the cold-start issue and further improves the recommendation performance. It is worth mentioning that the networks are trained with multi-task targets, e.g., CTR, CVR, etc. Following the previous online baseline, the method is implemented with the TensorFlow deep learning platform (tensorflow, ).
@@ A.4. Additional Test on Diverse Domains
We acknowledge that the applied domains of our proposed method are limited. We aim to test our proposed ODCRec on other domains, such as news recommendation and movie recommendation.
@@ A.4.1. Movie Recommendation
Specifically, for movie recommendation, we conducted experiments on the MovieLens 1M dataset (ML-1M). This dataset contains 1 million ratings from 6000 users on 4000 movies. In this experiment, we compared our proposed ODCRec with the baseline ICLRec. The experimental results are presented in the table below.
@@
From these experimental results, we can draw three conclusions. (1) Our proposed ODCRec achieves better recommendation performance, as evidenced by higher values for all four metrics: HR@5, HR@20, NDCG@5, and NDCG@20. For example, with the HR@5 metric, ODCRec outperforms ICLRec by 13.56%. (2) We calculated the $p$-value between our method and the baseline ICLRec. The results indicate that all the $p$-values are less than 0.05, suggesting that our ODCRec significantly outperforms ICLRec.
@@ A.5. Sensitivity
This section aims to answer research question (4). To verify the sensitivity of the proposed ODCRec, we test the performance on four datasets with different hyper-parameters. The experimental results are demonstrated in Figure 5. The x-axis denotes the values of hyper-parameters, and the y-axis denotes the values of the HR@5 metric. Concretely, for the trade hyper-parameter $\alpha$, we test the performance with $\alpha\in\{0.01,0.1,1,10,100\}$. We find that our proposed ODCRec is not very sensitive to trade-off $\alpha$. And ODCRec can achieve promising performance when $\alpha\in[0.1,10]$. Besides, for the cluster number $k$, we test the recommendation performance with $\alpha\in\{32,64,128,256,512\}$. The results show that ODCRec is also not very sensitive to cluster number $k$ and can perform well when $k\in[256,512]$.
@@
Figure 5. Sensitivity analyses of ODCRec. The results are evaluated by the HR@5 metric.
@@ A.7. Visualization
To further demonstrate ODCRec’s capability to capture users’ underlying intents, we conduct visualization experiments on four public datasets. Concretely, the learned behavior sequence embeddings are visualized via $t$-SNE during the training process. As shown in Figure 7, the first row to the fourth row denotes the results on Sports, Beauty, Toys, and Yelp, respectively. From these experimental results, we have three observations as follows. (1) At the beginning of training, the behavior sequence embeddings are disorganized and can not reveal the underlying intent of users. (2) During the training process, the latent distribution is optimized, and similar behaviors are grouped into the underlying intent. (3) After training, the users’ underlying intents appear, and we highlight them with circles in Figure 7. These intents can assist recommendation systems in better modeling users’ behavior and recommending items. In summary, the above experiments and observations verify the effectiveness of our proposed ODCRec.
@@ A.9. Solutions to Balance Problem
Balancing the different loss functions in our method is indeed an important challenge. Our overall loss function consists of next-item prediction loss, cluster-assisted contrastive loss, and cluster loss. It is formulated as follows: $\mathcal{L}_{\text{overall}}=\mathcal{L}_{\text{next\_item}}+0.1\times\mathcal{L}_{\text{ccl}}+\alpha\times\mathcal{L}_{\text{cluster}}$. We set the weight of $\mathcal{L}_{\text{ccl}}$ as 0.1 to maintain it in the same order of magnitude as the first term. This reduces the number of hyper-parameters and simplifies the selection process. The weight of $\mathcal{L}_{\text{cluster}}$ is set as a hyper-parameter $\alpha$. We test different values of $\alpha\in\{0.01,0.1,1,10,100\}$ and find that our ODCRec method is not very sensitive to the trade-off $\alpha$. Promising performance is achieved when $\alpha\in[0.1,10]$. The sensitivity analysis experiments are presented in Figure 5 (b) of the Appendix. In our proposed model, we set $\alpha$ to 1 for Sports and Toys datasets, 0.1 for the Yelp dataset, and 10 for the Beauty dataset. The selection of $\alpha$ is mainly based on the model performance. Balancing trade-offs in models with multiple loss functions or tasks is a challenge in many deep learning methods. There are various strategies to address this issue.
@@ A.11. Limitations & Future Work
In this paper, we propose a novel intent learning method named ODCRec based on the online differentiable clustering framework. It can better mine users’ underlying intents via unifying representation learning and clustering optimization. Besides, the online differentiable clustering module optimizes the clustering distribution via mini-batch data, thus improving the scalability and convenience of deployment. Moreover, we demonstrate the superiority, effectiveness, efficiency, sensitivity, convergence, and visualization of ODCRec on four benchmarks. ODCRec is also successfully applied in the real-world industrial application.
Although achieving promising results, we admit the proposed ODCRec algorithm has several limitations and drawbacks. We summarize them as follows.
  • Pre-defined Cluster Number. The cluster number in the ODCRec method is a pre-defined hyper-parameter. In the real-world data, it is hard to determine the cluster number, especially under the unsupervised conditions. In this paper, for the open benchmarks, we search the cluster number in 32, 64, 128, 256, 512. Besides, for the real-world application, the cluster number is set to 20 based on the number of user groups. However, either the search method or the expert knowledge can not determine the cluster number well at once. The cluster number may change dynamically during model training, and the proposed method may fail to achieve promising performance.
  • Limited Recommendation Domains. In this paper, we adopt four recommendation datasets, including Sports, Beauty, Toys, and Yelp. But, these four datasets are all business recommendation datasets. Besides, in the industrial application aspect, the used data is also a business recommendation dataset. Therefore, the recommendation domain is limited and this paper merely demonstrates the superiority of ODCRec in one domain.
  • No Strategy for Solving Rapid Preference Change. In real-world applications, the intents of users usually change rapidly. Although the proposed ODCRec can capture the user’s intent effectively in the training data, the user’s intent may change dynamically in the future. However, this work does not provide any strategies to alleviate this issue, which may lead to unpromising performance in real-world applications.
@@
  • More Recommendation Domains. As mentioned above, we admit applied domains of our proposed method are limited. We aim to test our proposed ODCRec on other domains like news recommendation and movie recommendation in the future. For the news recommendation, we can conduct experiments on MIND dataset. This dataset is a large-scale dataset for news recommendation research. It was collected from anonymized behavior logs of the Microsoft News website. MIND contains about 160k English news articles and more than 15 million impression logs generated by 1 million users. For the movie recommendation, we conduct experiments on the MovieLens 1M dataset (ML-1M), Netflix, Douban, Twitch-100k, Twith-full datasets, etc.
  • More Recommendation Domains. As mentioned above, in the real-world application, the intents of the users may change rapidly. However, there is no strategy in the proposed ODCRec method to solve this problem. To this end, in the future, we can propose new cluster center update strategies, such as momentum updates, to control the change rate of the users’ intents. Namely, $\mathbf{C}_{t}=m\cdot\mathbf{C}_{t}+(1-m)\cdot\mathbf{C}_{t-1}$. Here, $\mathbf{C}_{t}$ denote the cluster center embeddings at $t$ and $m$ denotes the momentum. Then, the cluster centers (intents of users) will be changed rapidly when $m$ is large, and the cluster centers (intents of users) will be changed slowly when $m$ is small. This strategy will control the change rate of the users’ intent embeddings, therefore alleviating the above problem.